Star Company’s accountants estimate total overhead for each month will be $64,000. They will allocate overhead on the basis of direct labour cost. During the current month, three jobs were worked on: Job 745 Job 746 Job 747 Direct material $36,000 $56,000 $24,000 Direct labour $56,000 $72,000 $40,000 Job 745 was completed and sold, Job 746 was completed and Job 747 is still in process. Budgeted direct labour cost for the month was $160,000 and actual overhead was $65,800. Assuming overhead was $100 underapplied and that this amount is considered immaterial, the final adjusted balance of cost of goods sold is:

$92,100

$211,900

$114,500

$79,900
Answer #1

Only Job 745 is sold. Cost of goods sold arises only for Job 745.

Overhead rate = 64,000 / 160,000 = 40%

Total cost of Job 745 = $36,000 + $56,000 + ($56,000 X 40%) + $100 = $114,500

3rd option.